Determine the points where the curve 2x+4y^2-y=1has a vertical tangent line
1
Expert's answer
2020-10-12T18:54:14-0400

The given equation of curve is

"2x+4y^2-y=1"

Since, vertical tangent found to those points where slope="dy\/dx" is "\pm \infty"

Now, take derivative both sides of the above curve with respect to x we get


"2+8y\frac{dy}{dx}-\frac{dy}{dx}=0\\
\implies \frac{dy}{dx}=\frac{2}{1-8y}"

Thus,"\frac{dy}{dx}=\pm\infty" only at "y=1\/8" , hence on plugin this value to the equation of curve we get,


"2x+4(1/8)^2-(1/8)=1\\
\implies 2x=\frac{17}{16}\implies x=\frac{17}{32}"

Thus, point where the curve "2x+4y^2-y=1" has vertical tangent is "(17\/32,1\/8)"